OpenStudy (anonymous):

i need help simplifing m to the negative 5 power over m to the 6 power

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[m ^{-5}/m ^{6}\]

OpenStudy (anonymous):

if you are dividing exponents with the same base (m) then all you have to do is subtract the top exponent from the bottom. so -5-6=-11. m^-11

OpenStudy (anonymous):

sorry-- bottom exponent from the top

OpenStudy (anonymous):

so its\[1/m ^{11}\] right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

m*-5=1/m*5 (1/m*5)/m*6=(1/m*5)x(1/m*6)=1/m*11